OpenStudy (anonymous):

Can someone help me with this? In 1992, the life expectancy of males in a certain country was 62.5 years. In 1997, it was 66.3 years. Let E represent the life expectancy in year t and let t represent the number of years since 1992. E(t)=?t+? Use the function to predict the life expectancy of males in 2006 E(14)=?

OpenStudy (saifoo.khan):

Online college/

OpenStudy (anonymous):

I know right

OpenStudy (anonymous):

The college sucks because they give you equations that were not covered in the material and expect you to figure them out

OpenStudy (saifoo.khan):

hmm, ikr.. can i have experienced this 3 times.

OpenStudy (anonymous):

well never go to Kaplan

OpenStudy (saifoo.khan):

Dang it.. u in Kaplan too!

OpenStudy (anonymous):

yea

OpenStudy (saifoo.khan):

i got a student, she is in Kaplan too! lol

OpenStudy (anonymous):

yea, the university really sucks....no help whatsoever

OpenStudy (saifoo.khan):

E(t)=0.76t+62.5

OpenStudy (saifoo.khan):

E(t)=0.76(14)+62.5 E(t)=73.14

OpenStudy (anonymous):

what if both need to be rounded to the nearest tenth?

OpenStudy (saifoo.khan):

they are.

OpenStudy (anonymous):

oh, see how dumb i am

OpenStudy (anonymous):

i should have been blonde

OpenStudy (saifoo.khan):

lol, all college students have the same problem! :(

OpenStudy (saifoo.khan):

Lol.

OpenStudy (anonymous):

are you sure about the tenths? someone is telling me that 62.5 rounded to the nearest tenth is 60

OpenStudy (saifoo.khan):

Sorrry mybad this time! :( E(t)=0.8t+62.5 E(t)=0.8(14)+62.5 E(t)=73.7

OpenStudy (saifoo.khan):

No.

OpenStudy (anonymous):

73.1?

OpenStudy (saifoo.khan):

73.7

OpenStudy (anonymous):

Im sorry, I have someone arguing with me on this end

OpenStudy (saifoo.khan):

lol, NP

OpenStudy (anonymous):

they are saying that 73.14 rounded to the nearest tenth is 73.1

OpenStudy (saifoo.khan):

yes

OpenStudy (saifoo.khan):

they are correct.

OpenStudy (anonymous):

wait a minute, so its not 73.7?

OpenStudy (saifoo.khan):

why?

OpenStudy (anonymous):

how did the second answer come to 73.7?

OpenStudy (saifoo.khan):

coz, 0.73 is now changed to 0.76 is now 0.8... after rounding it off.